# Pietro Cassieri+, Gianluca Fimiani+, Simone Romano+, Marina Garofano+, Alessia Bramanti+, Giuseppe Scanniello+

### v1

@@ Abstract
The management of chronic heart failure (CHF) presents significant challenges in modern healthcare, requiring continuous monitoring, early detection of exacerbations, and personalized treatment strategies. This paper presents the PrediHealth research project and its preliminary results in the context of heart failure (HF). PrediHealth, conducted in collaboration with a software industry operating in the CPS/IoT fields, aims to address the challenges above by integrating telemedicine, mobile health (m-health) solutions, and predictive analytics into a unified digital healthcare platform. We leveraged an interoperable IoT web-based application, a telemonitoring kit equipped with medical devices and environmental sensors, and AI-driven predictive models to support clinical decision-making. The project follows a structured methodology comprising research on emerging CPS/IoT technologies, system prototyping, predictive model development, and empirical validation within the healthcare ecosystem. Through its innovative AI-enhanced approach, PrediHealth improves patient engagement, reduces hospitalization rates, enhances continuity of care, and lowers healthcare costs.
@@ I Introduction
Cardiovascular diseases (CVDs) remain the leading cause of mortality, morbidity, hospitalization, and disability in many countries, imposing a significant burden on healthcare systems. For example, direct healthcare costs related to CVDs account in Italy for approximately 65% of all public health expenditures, with pharmaceutical expenses constituting an additional 25%. Heart failure (HF) is particularly concerning as it is one of the fastest-growing cardiovascular conditions, primarily due to the increasing life expectancy of the population. The Diagnosis-Related Group (DRG) 127, which pertains to HF, ranked as the second most frequent DRG in Italy, with 223,240 hospitalizations reported in 2015. Furthermore, within 30 days of hospital discharge, the mortality rate for HF patients stands at 11%, with readmission and rehospitalization rates of 14% and 3%, respectively [1]. Things are similar in different countries, where HF represents a significant healthcare challenge. In Germany, for example, HF has become the most common reason for hospitalization. Notably, in 2017, HF was the leading cause of in-hospital death, accounting for 8.9% of all cases, with a higher rate in East Germany (65 deaths per 100,000 population) compared to West Germany (43 deaths per 100,000 population) [2]. In the United Kingdom, HF is responsible for 1–2% of total healthcare expenditure, reflecting its substantial economic impact [3]. In the United States, approximately 6 million individuals were living with HF between 2015 and 2018, with prevalence projected to increase by 46% between 2012 and 2030 [4].
@@
Website: https://docenti.unisa.it/020007/ricerca/progetti?progetto=62586
@@
Industrial Partner: MediNet S.r.L.
@@
In this paper, we present the PrediHealth research project (see Table I for some additional information). This project focuses on using CPS (cyber-physical systems) and IoT (Internet of Things) solutions [^1] While CPS encompasses a broader range of networked systems, IoT specifically focuses on internet-enabled connectivity. in smart healthcare and medical research. The project is conducted in collaboration with a software industrial partner operating in CPS/IoT fields and aims to facilitate the integrated hospital-territory-home management of HF patients. It leverages state-of-the-art technologies such as telemedicine, CPS/IoT devices (wearable and not), artificial intelligence (AI), and machine learning (ML).
@@
  • A web-based IoT platform to securely use and exchange telemonitoring data and ensure integration with existing hospital electronic health records. To achieve this, we conducted a comprehensive review of open-source IoT-based web applications, evaluating them against a set of predefined criteria, including security, interoperability, scalability, and ease of integration. Based on this analysis, we selected the most suitable application and adapted it to meet the specific requirements of our healthcare context.
  • A telemonitoring kit, comprising medical-grade wearable devices and environmental sensors, designed for real-time health tracking of HF patients. To ensure its effectiveness and suitability, we conducted a comprehensive review of these devices/sensors, assessing them against a set of predefined criteria, including accuracy, reliability, interoperability, and patient comfort.
  • A decision-making support system, developed using AI/ML-powered algorithms to analyze patient-centered data, facilitating patients’ stratification [^2] It consists in categorizing individuals into different risk groups based on their health status, disease progression, or other factors. and early intervention. The system is designed to identify high-risk patients who would benefit from continuous remote monitoring. The final goal is to determine which patients require a telemonitoring kit to be proactively enrolled in a personalized monitoring program.
@@ II Background, Related Work, and Motivation
In this section, we review related work and background information relevant to the PrediHealth project and its foundational research, with a focus on (i) remote patient monitoring, (ii) mobile health (m-health) solutions, and (iii) predictive models for cardiovascular diseases.
@@
Remote Patient Monitoring (RPM) and Telemonitoring: RPM and telemonitoring are closely related concepts. Specifically, RPM is a broad term that refers to the use of digital technologies to collect health data from patients outside traditional healthcare settings (e.g., at home) and transmit it to healthcare providers. On the other hand, telemonitoring is a subset of RPM that specifically focuses on continuous or periodic remote monitoring of patients with chronic diseases, such as HF. Telemonitoring emphasizes real-time or near-real-time data collection, trend analysis, and early detection of worsening conditions to enable timely interventions. In short, RPM is a broader category that includes various remote data collection methods, while telemonitoring is more focused on continuous disease-specific monitoring and proactive healthcare interventions. Studies have demonstrated the effectiveness of RPM/telemonitoring in improving patient outcomes, particularly for chronic conditions like HF. In particular, Bashshur et al. [6] conducted a systematic review examining telemedicine’s impact on chronic diseases, including HF, stroke, and Chronic obstructive pulmonary disease. Their findings reveal reductions in mortality (by 15% to 56% in HF patients), hospital admissions, and healthcare costs. Similarly, YM et al. [7] highlights the role of telemedicine consultation and telemonitoring in managing hypertension, diabetes, and rheumatoid arthritis. The study emphasizes the importance of long-term interventions.
@@
Mobile Health (m-health): m-health and telemonitoring are closely linked, as both leverage digital technologies to enhance remote patient care. M-health primarily focuses on mobile applications and wearable devices that enable real-time health tracking, patient engagement, and self-management. Telemonitoring extends this concept by integrating m-health data with broader healthcare systems. M-health applications have shown considerable promise in enhancing patient engagement and promoting self-management, particularly for individuals with chronic diseases such as heart failure (HF). The work of Stampe et al. [8] explored the use of m-health technologies in chronic disease management, emphasizing their role in improving patient-physician consultations. Researchers also observed that the use of these technologies helped in better decision-making by both patients and clinicians. Schmaderer et al. [9] conducted a pilot study focused on m-health self-management interventions for patients with heart failure. Their results indicated significant improvements in symptom management and overall health outcomes among patients who utilized the mobile health app. Specifically, the study found a notable reduction in hospitalizations and readmissions, suggesting that m-health applications could be instrumental in managing HF symptoms and preventing exacerbations.
@@
Predictive Models: Predictive AI/ML-powered models play a crucial role in managing chronic diseases like HF by enabling early detection, risk stratification, and personalized treatment for patients. These models analyze large datasets (e.g., from wearable and environmental sensors) to support proactive interventions, reduce hospitalizations, and improve patient outcomes, making predictive modeling a key component of modern telemedicine and, more specifically, telemonitoring. For example, Medhi et al. [10] conducted a review by selecting 150 studies from databases such as PubMed, Google Scholar, and the Cochrane Library. Their synthesis detailed how AI/ML-powered models—spanning classical methods like SVMs, Random Forest, and decision trees to advanced deep learning approaches—can process large datasets from wearables, imaging, and other clinical sources to enable early detection, risk stratification, and personalized treatment of heart failure. Their work highlights the potential of these predictive models to reduce hospitalizations and improve patient outcomes and emphasizes the transformative role of AI in modern telemedicine and telemonitoring. Visco et al. [11] developed a predictive model for cardiovascular diagnostics. Their study identified correlations between worsening HF and three key parameters: Creatinine, sPAP, and CAD. To enhance interpretability, they introduced a GP-based classifier that achieved 97% accuracy, outperforming traditional ML models.
@@
Summary: Existing literature supports the potential impact of integrating telemonitoring, m-health, and predictive models in managing chronic conditions like HF. The PrediHealth project builds on these insights by providing an integrated solution to improve the personalization of care, enhance patient engagement, and prevent exacerbations. Furthermore, the project contributes to the growing body of knowledge on the use of AI/ML models in healthcare, with a focus on HF.
@@ III Objectives
The high-level workflow of PrediHealth begins with the stratification of patients at risk of HF using AI/ML-powered predictive models. Patients identified as high-risk are then equipped with wearable and environmental sensors, which are seamlessly integrated into a telemonitoring system based on an IoT-enabled web platform. In PrediHealth, we defined the following objectives:
@@
O1: Develop an interoperable web-based IoT platform. The objective is to develop an interoperable web-based IoT platform. The platform has to establish a secure, scalable, and standardized digital infrastructure for the seamless collection, transmission, and analysis of health data from telemonitoring devices. This platform facilitates real-time monitoring of HF patients by integrating wearable medical devices and environmental sensors, ensuring continuous and remote patient monitoring. Additionally, the platform has to incorporate advanced security mechanisms, including encryption and role-based access controls, to protect sensitive patient data and comply with relevant regulatory frameworks (e.g., GDPR). The web-based architecture will also support scalability and flexibility, enabling future integration with different telemonitoring devices and services. A key aspect of O1 is to have a platform that supports HL7 FHIR (Fast Healthcare Interoperability Resources), a widely accepted healthcare data exchange standard that enables interoperability across diverse healthcare ecosystems. By adhering to this standard, the PrediHealth platform ensures compatibility with hospital information systems, electronic health records, and other telemedicine solutions.
@@
O2: Design a telemonitoring kit for HF patients. This objective focuses on the development of a comprehensive telemonitoring kit specifically designed for patients with HF, integrating advanced (wearable and not) medical devices and environmental sensors to enable continuous health monitoring. The kit ensures non-intrusive, user-friendly, and clinically effective monitoring. Key components of the kit include wearable medical devices, such as smartwatches or biosensors, capable of tracking critical physiological parameters like heart rate, heart rate variability (HRV), blood pressure, oxygen saturation (SpO2), and electrocardiographic (ECG) signals. Additionally, environmental sensors will monitor factors such as air quality, temperature, and humidity since these elements could affect cardiovascular health. The devices of the telemonitoring kit will be interoperable with the broader PrediHealth platform (i.e., the platform to be realized in O1), ensuring secure and standardized data transmission.
@@
O3: Develop a multimarker score to support decision-making. The goal of O3 is twofold: (i) building AI/ML-powered models to facilitate the risk stratification of patients vulnerable to acute exacerbations or disease progression for HF and (ii) defining thresholds and scores for risk factors associated with HF exacerbations of patients when using telemonitoring kits. As for the first point, a decision support system (DSS) developed using advanced AI/ML algorithms is designed to analyze patient-centered data and facilitate risk stratification. This system leverages historical medical records to generate a personalized risk for each patient. The primary function of this DSS is to identify high-risk patients who are most vulnerable to acute exacerbations or disease progression for HF. This allows making informed decisions about the patient—determining who would benefit most from a personalized telemonitoring program—and then allocating such patient to that program. As for the second point, we aim to design and implement a multimarker score that integrates multiple parameters (both personal and environmental) to enhance clinical decision-making in HF management. This score takes into account real-time data from telemonitoring devices: IoT-enabled medical devices and environmental sensors. The multimarker score is developed by identifying key biomarkers and risk factors associated with HF exacerbations, such as heart rate variability, oxygen saturation, respiratory rate, blood pressure fluctuations, body weight changes, and environmental conditions (e.g., air quality, humidity, and temperature). This aims to detect early signs of decompensation, enabling timely interventions and reducing the likelihood of emergency hospitalizations. In case the patient is not at home, only the data from IoT-enabled medical devices will contribute to the score to support decision-making.
@@
O4: Demonstrate and validate the proposed prototype. We empirically evaluate the research solutions implemented in PrediHealth to ensure its technical feasibility, clinical relevance, and usability within real-world healthcare environments. The validation will take place within THE (Tuscany Health Ecosystem), [^3] https://tuscanyhealthecosystem.it/en/ a collaborative framework that integrates healthcare providers, research institutions, and technology developers. The evaluation encompasses both clinical and technical assessments. The clinical validation will involve testing the predictive accuracy of AI/ML models, ensuring that risk stratification and scores to support decision-making provide clinically meaningful insights. The technical validation will focus on evaluating the platform and the telemonitoring kit in terms of data acquisition, transmission, usability, and integration with healthcare records. The findings from the empirical evaluation form the foundation for potential large-scale clinical trials, paving the way for broader adoption of the PrediHealth solutions. s
@@ IV Methodology and Structure
Given the objectives mentioned above, the project is organized into four Work Packages (WPs). Further details on the WPs are reported in Table II. Specifically, WP1, led by the Department of Computer Science at the University of Salerno (UniSA DI), addresses O5 and focuses on project coordination and dissemination of results throughout the entire project duration. WP2, coordinated by the University of Salento (UniSalento), covers the first five months and focuses on studying emerging technologies with potential healthcare applications at the international level. To some extent, this WP is cross-cutting to O1, O2, and O3 because each of the objectives necessitated a specific examination of the state of the art and practice. WP3, also led by UniSalento, addresses O1 and O2 and begins in the fourth month and is dedicated to deploying a web-based IoT platform and designing a telemonitoring kit for HF patients. Finally, WP4, managed by the Department of Medicine at the University of Salerno (UniSA DipMed), addresses O3 and O4 and is centered around data collection, the training of the predictive model, and the definition of scores for risk factors associated with HF exacerbations. The implemented research solutions are empirically evaluated in WP3 and WP4 and then both these WPs address O4. The industrial partner of PrediHealth contributed to the implementation aspects of WP3 and WP4, i.e., the technical WPs of the project. A description of the project WPs follows.
WP1: Coordination and dissemination of results. It spans the full project duration and includes the definition and implementation of the project management model and the definition of strategies for the dissemination of the scientific results produced in the project (see O5).
WP2: Study of emerging technologies and solutions. Its goal is to scout innovative and advanced medical devices and environmental sensors for the telemonitoring kit for HF patients. Additionally, it involves analyzing telemedicine platforms and web-based IoT platforms, technological solutions for integrated chronic patient management, the critical issues of current SC monitoring platforms, and AI/ML-powered predictive models for the early detection of HF diseases and risk stratification patients. WP2 is cross-cutting to O1, O2, and O3 because the systematic study of the state of the art and practice is vital for the PrediHealth success.
@@
WP4: Definition and training of the predictive model. This WP aims to study and define AI/ML-powered models to facilitate the risk stratification of patients vulnerable to acute exacerbations or disease progression for HF. Our primary goal is to facilitate the utilization of publicly available datasets to train AI/ML-powered models, ensuring robust and data-driven development within the project. In case publicly available datasets are not available, we plan to use datasets of the THE ecosystem. Additionally, WP4 focuses on the prototypical development and assessment of predictive models and algorithms. WP4 also aims to define thresholds and scores for HF risk factors when using the telemonitoring kit.

### v2

# Alessia Bramanti+, Giuseppe Scanniello+

@@ Abstract
The management of chronic heart failure presents significant challenges in modern healthcare, requiring continuous monitoring, early detection of exacerbations, and personalized treatment strategies. This paper presents the preliminary results of the PrediHealth research project conducted in this context. Specifically, it aims to address the challenges above by integrating telemedicine, mobile health solutions, and predictive analytics into a unified digital healthcare platform. We leveraged a web-based IoT platform, a telemonitoring kit with medical devices and environmental sensors, and AI-driven predictive models to support clinical decision-making. The project follows a structured methodology comprising research on emerging CPS/IoT technologies, system prototyping, predictive model development, and empirical validation.
@@ I Introduction
CardioVascular Diseases (CVDs) remain the leading cause of mortality, morbidity, hospitalization, and disability in many countries, imposing a significant burden on healthcare systems. For example, direct healthcare costs related to CVDs account for approximately 65% of all public health expenditures in Italy, with pharmaceutical expenses constituting an additional 25%. Heart Failure (HF) is particularly concerning as it is one of the fastest-growing cardiovascular conditions, primarily due to the increasing life expectancy of the population. The Diagnosis-Related Group (DRG) 127, which pertains to HF, ranked as the second most frequent DRG in Italy, with 223,240 hospitalizations reported in 2015. Furthermore, within 30 days of hospital discharge, the mortality rate for HF patients stands at 11%, with readmission and rehospitalization rates of 14% and 3%, respectively [1]. Things are similar in different countries, where HF represents a significant healthcare challenge. In Germany, for example, HF has become the most common reason for hospitalization. Notably, in 2017, HF was the leading cause of in-hospital death, accounting for 8.9% of all cases, with a higher rate in East Germany (65 deaths per 100,000 population) compared to West Germany (43 deaths per 100,000 population) [2]. In the United Kingdom, HF is responsible for 1–2% of total healthcare expenditure, reflecting its substantial economic impact [3]. In the United States, approximately 6 million individuals were living with HF between 2015 and 2018, with prevalence projected to increase by 46% between 2012 and 2030 [4].
@@
Website: https://tinyurl.com/2mwbu7r4
@@
Industrial Partner: MediNet S.r.L. (https://www.medinetsrl.eu)
@@
In this paper, we present the PrediHealth research project (see Table I for some additional information). This project focuses on using CPS (Cyber-Physical Systems) and IoT solutions [^1] While CPS encompasses a broader range of networked systems, IoT specifically focuses on internet-enabled connectivity. in smart healthcare and medical research. The project is conducted in collaboration with a software industrial partner operating in CPS and IoT fields and aims to facilitate the integrated hospital-territory-home management of HF patients. It leverages state-of-the-art technologies such as telemedicine, CPS/IoT devices (wearable and not), Artificial Intelligence (AI), and Machine Learning (ML).
@@
  • A web-based IoT platform to securely use and exchange telemonitoring data and ensure integration with existing hospital electronic health records. To achieve this, we conducted a review of open-source IoT-based web applications, evaluating them against a set of criteria, including security, interoperability, scalability, and ease of integration. Based on this analysis, we selected the most suitable application and adapted it to meet the specific requirements of our project.
  • A telemonitoring kit, comprising medical-grade wearable devices and environmental sensors, designed for real-time health tracking of HF patients. To ensure its effectiveness and suitability, we conducted a review of these devices/sensors, assessing them against a set of predefined criteria, including accuracy, reliability, interoperability, and patient comfort.
  • A decision-making support system, developed using AI/ML-powered algorithms to analyze patient-centered data, facilitating patients’ stratification [^2] It consists in categorizing individuals into different risk groups based on their health status, disease progression, or other factors. and early intervention. The system is designed to identify HF risk patients who would benefit from continuous remote monitoring. The final goal is to determine which patients require a telemonitoring kit to be proactively enrolled in a personalized monitoring program.
@@ II Background, Related Work, and Motivation
In this section, we review related work and background information relevant to the PrediHealth project and its foundational research, with a focus on (i) Remote Patient Monitoring (RPM) and telemonitoring, (ii) mobile health (m-health) solutions, and (iii) predictive models for CVDs.
@@
RPM and Telemonitoring. RPM and telemonitoring are closely related concepts. Specifically, RPM is a broad term that refers to the use of digital technologies to collect health data from patients outside traditional healthcare settings (e.g., at home) and transmit it to healthcare providers. On the other hand, telemonitoring is a subset of RPM that specifically focuses on continuous or periodic remote monitoring of patients with chronic diseases, such as HF. Telemonitoring emphasizes real-time or near-real-time data collection, trend analysis, and early detection of worsening conditions to enable timely interventions. In short, RPM is a broader category that includes various remote data collection methods, while telemonitoring is more focused on continuous disease-specific monitoring and proactive healthcare interventions. Studies have demonstrated the effectiveness of RPM/telemonitoring in improving patient outcomes, particularly for chronic conditions like HF. In particular, Bashshur et al. [6] conducted a systematic review examining telemedicine’s impact on chronic diseases, including HF, stroke, and chronic obstructive pulmonary disease. Their findings reveal reductions in mortality (by 15% to 56% in HF patients), hospital admissions, and healthcare costs. Similarly, Ma et al. [7] highlight the role of telemedicine consultation and telemonitoring in managing hypertension, diabetes, and rheumatoid arthritis. The study emphasizes the importance of long-term interventions.
@@
M-health. m-health and telemonitoring are closely linked, as both leverage digital technologies to enhance remote patient care. M-health primarily focuses on mobile applications and wearable devices that enable real-time health tracking, patient engagement, and self-management. Telemonitoring extends this concept by integrating m-health data with broader healthcare systems. M-health applications have shown considerable promise in enhancing patient engagement and promoting self-management, particularly for individuals with chronic diseases such as HF. Stampe et al. [8] explored the use of m-health technologies in chronic disease management, emphasizing their role in improving patient-physician consultations. The researchers also observed that the use of these technologies helped in better decision-making by both patients and clinicians. Schmaderer et al. [9] conducted a pilot study focused on m-health self-management interventions for patients with HF. Their results indicated significant improvements in symptom management and overall health outcomes among patients who utilized m-health apps. Specifically, the study found a notable reduction in hospitalizations and readmissions, suggesting that m-health apps could be instrumental in managing HF symptoms and preventing exacerbations.
@@
Predictive Models for CVDs. Predictive AI/ML-powered models play a crucial role in managing chronic diseases like HF by enabling early detection, risk stratification, and personalized treatment for patients. These models analyze large datasets (e.g., from wearable and environmental sensors) to support proactive interventions, reduce hospitalizations, and improve patient outcomes, making predictive modeling a key component of modern telemedicine and, more specifically, telemonitoring. For example, Medhi et al. [10] conducted a review by selecting 150 studies from databases such as PubMed, Google Scholar, and Cochrane Library. Their synthesis detailed how AI/ML-powered models—spanning classical methods like support vector machines, random forest, and decision trees to advanced deep learning approaches—can process large datasets from wearables, imaging, and other clinical sources to enable early detection, risk stratification, and personalized treatment of HF. Their work highlights the potential of these predictive models to reduce hospitalizations and improve patient outcomes and emphasizes the transformative role of AI in modern telemedicine and telemonitoring. Visco et al. [11] developed a predictive model for cardiovascular diagnostics. Their study identified correlations between worsening HF and three key parameters: creatinine, systolic pulmonary artery pressure, and coronary artery disease. To enhance interpretability, they introduced a classifier based on genetic programming that achieved 97% accuracy, outperforming traditional ML models.
@@
Summary. Existing literature supports the potential impact of integrating telemonitoring, m-health, and predictive models in managing chronic conditions like HF. The PrediHealth project builds on these insights by providing an integrated solution to improve the personalization of care, enhance patient engagement, and prevent exacerbations. Furthermore, the project contributes to the growing body of knowledge on the use of AI/ML models in healthcare, with a focus on HF.
@@ III Objectives
The high-level workflow of the PrediHealth project begins with the stratification of patients at risk of HF using an AI/ML-powered predictive model. Patients identified as at-risk are then equipped with wearable and environmental sensors, which are integrated into a telemonitoring system based on a web-based IoT platform. To achieve the overarching goal of PrediHealth, we have then defined and pursued the following operational objectives:
@@
O1: Develop an interoperable web-based IoT platform. The objective is to develop an interoperable web-based IoT platform. The platform has to establish a secure, scalable, and standardized digital infrastructure for the collection, transmission, and analysis of health data from telemonitoring devices. This platform facilitates real-time monitoring of HF patients by integrating wearable and non-wearable medical devices as well as environmental sensors. Additionally, the platform has to incorporate advanced security mechanisms, including encryption and role-based access controls, to protect sensitive patient data and comply with relevant regulatory frameworks (e.g., GDPR). The web-based architecture will also support scalability and flexibility, enabling future integration with different telemonitoring devices and services. A key aspect of O1 is to have a platform that supports HL7 FHIR (Fast Healthcare Interoperability Resources), a healthcare data exchange standard that enables interoperability across diverse (healthcare) ecosystems. By adhering to this standard, the platform developed in PrediHealth ensures compatibility with hospital information systems, electronic health records, and other telemedicine solutions.
@@
O2: Develop a telemonitoring kit for HF patients. This objective focuses on the development of a comprehensive telemonitoring kit specifically designed for patients with HF, integrating advanced (wearable and not) medical devices and environmental sensors to enable continuous health monitoring. The kit ensures non-intrusive, user-friendly, and clinically effective monitoring. Key components of the kit include wearable medical devices, such as smartwatches and scales, capable of tracking critical physiological parameters like, for example, heart rate, weight, Heart Rate Variability (HRV), blood pressure, oxygen saturation (SpO2), and ElectroCardioGraphic (ECG) signals. Additionally, environmental sensors will monitor factors such as air quality, temperature, and humidity since these elements could affect cardiovascular health. The devices of the telemonitoring kit will be interoperable with the platform developed in PrediHealth.
@@
O3: Develop a multimarker score to support decision-making. The goal of O3 is twofold: (i) building AI/ML-powered models to facilitate the risk stratification of patients vulnerable to acute exacerbations or disease progression for HF (i.e., identify patients at HF risk) and (ii) defining thresholds and scores for risk factors associated with HF exacerbations of patients (those previously considered at HF risk) when using telemonitoring kits. As for the first point, a Decision Support System (DSS) developed using advanced AI/ML algorithms is designed to analyze patient-centered data and facilitate risk stratification. This system leverages historical medical records to generate an HF risk for each patient. That is, the primary function of this DSS is to identify HF risk patients who are most vulnerable to acute exacerbations or disease progression for HF. This allows making informed decisions about the patient–determining who would benefit the most from a telemonitoring program–and then allocating such a patient to that program. As for the second point, we aim to design and implement a multimarker score that integrates multiple parameters (both personal and environmental) to enhance clinical decision-making in HF management. This score takes into account real-time data from telemonitoring devices: IoT-enabled (wearable and not) medical devices and environmental sensors. The multimarker score is developed by identifying key biomarkers and risk factors associated with HF exacerbations, such as HRV, SpO2, respiratory rate, blood pressure fluctuations, body weight changes, and environmental conditions (e.g., air quality, humidity, and temperature). This aims to detect early signs of decompensation, enabling timely interventions and reducing the likelihood of emergency hospitalizations. In case the patient is not at home, only the data from IoT-enabled medical devices will contribute to the score to support decision-making.
@@
O4: Demonstrate and validate the proposed prototype. We empirically evaluate the research solutions implemented in PrediHealth to ensure its technical feasibility, clinical relevance, and usability within real-world healthcare environments. Our validation will take place within THE (Tuscany Health Ecosystem) [12] a collaborative framework that integrates healthcare providers, research institutions, and technology developers. The evaluation encompasses both clinical and technical assessments. The clinical validation will involve testing the predictive accuracy of AI/ML models, ensuring that risk stratification and scores to support decision-making provide clinically meaningful insights. The technical validation will focus on evaluating the platform and the telemonitoring kit in terms of data acquisition, transmission, usability, and integration with healthcare records. The outcomes from the empirical evaluation form the foundation for potential large-scale clinical trials, paving the way for broader adoption of the PrediHealth solutions.
@@ IV Methodology and Structure
Given the operational objectives above, the project is organized into four Work Packages (WPs). Details on the WPs are reported in Table II. Specifically, WP1, led by the Department of Computer Science at the University of Salerno (UniSA DI), addresses O5 and focuses on project coordination and dissemination of results throughout the entire project duration. WP2, coordinated by the University of Salento (UniSalento), covers the first five months and focuses on studying emerging technologies with potential healthcare applications at the international level. To some extent, this WP is cross-cutting to O1, O2, and O3 because each of these objectives required a specific examination of the state of the art and practice. WP3, also led by UniSalento, addresses O1 and O2. It begins in the fourth month and is dedicated to the deployment of a web-based IoT platform and a telemonitoring kit for HF patients. Finally, WP4, managed by the Department of Medicine at the University of Salerno (UniSA DipMed), addresses O3 and O4. It is centered around the data collection, the training of the predictive model, and the definition of scores for HF risk factors. The implemented research solutions are empirically evaluated in WP3 and WP4, and then both these WPs address O4. The industrial partner of PrediHealth contributed to the implementation aspects of WP3 and WP4, i.e., the technical WPs of the project. A description of the project WPs follows.
WP1: Coordination and dissemination of results. It spans the full project duration and includes the definition and implementation of the project management model and the definition of strategies for the dissemination of the scientific results produced in the project.
WP2: Study of emerging technologies and solutions. Its goal is to scout innovative and advanced medical devices and environmental sensors for the telemonitoring kit for HF patients. Additionally, it involves analyzing telemedicine platforms and web-based IoT platforms, technological solutions for integrated chronic patient management, the critical issues of current SC monitoring platforms, and AI/ML-powered predictive models for the early detection of HF diseases and risk stratification patients.
@@
WP4: Definition and training of the predictive model. This WP aims to study and define AI/ML-powered models to facilitate the stratification of patients at HF risk. Our primary goal was to facilitate the utilization of publicly available datasets (if any) to train AI/ML-powered models, ensuring robust and data-driven development within the project. As an alternative, we planned to use datasets of the THE ecosystem. Additionally, WP4 focuses on the prototypical development and assessment of predictive models and algorithms. WP4 also aims to define thresholds and scores for HF risk factors when using the telemonitoring kit.